# Whitehead’s problem and condensed mathematics

### v3

@@ 4. Countable groups
We can thus injectively enumerate $B_{K}$ as $\langle t_{n}\mid n<\omega\rangle$ and, for each $n<\omega$, let $K_{n}$ denote the subgroup of $K$ generated by $\{t_{k}\mid k<n\}$. We will define our homomorphism $\varphi:K\rightarrow\mathbb{Z}$ by recursion, requiring that $\varphi(t_{n})\in\{0,1\}$ for all $n<\omega$. We will do this by constructing a strictly increasing sequence $\langle n_{i}\mid i<\omega\rangle$ of natural numbers and, for each $i<\omega$, a homomorphism $\varphi_{i}:K_{n_{i}}\rightarrow\mathbb{Z}$ in such a way that $\varphi_{i}[\{t_{k}\mid k<n_{i}\}]\subseteq\{0,1\}$ and, for all $i<j<\omega$, $\varphi_{j}\restriction K_{n_{i}}=\varphi_{i}$.
@@
Then $s_{i}^{*}\in K$ is of the form $j_{i}z_{i}-r$, where $r\in F_{i}$. Moreover, since all of the coefficients $b_{k}$ for $k\in[n_{i},n_{i+1})$ are divisible by $j_{i}$, it follows that $s_{i}^{*}$ is divisible by $j_{i}$ (in $K$). Since $r=j_{i}z_{i}-s_{i}^{*}$, this means that there exists $r^{*}\in F_{i}$ such that $r=j_{i}r^{*}$. We get $s_{i}^{*}=j_{i}(z_{i}-r^{*})\in K$ which implies $z_{i}-r^{*}\in K$ using the fact that $A$ is torsion-free. In particular, $\pi(z_{i})=\pi(r^{*})$, so $\pi(z_{i})\in\pi[F_{i}]$, contradicting the fact that $\pi(z_{i})=x_{i}\notin\pi[F_{i}]$. ∎
@@
In the formulation of Theorem 4.2, we could not have hoped to find $\varphi$ with $(\varphi\restriction B_{K})^{-1}\{1\}$ finite, let alone a $\varphi$ satisfying the stronger condition (2) from Proposition 4.4. The reason is that any given countable torsion-free group $A$, being the union of a countable chain of finitely generated free subgroups, is isomorphic to a (pure) subgroup of $\mathbb{Z}^{\omega}/\mathbb{Z}^{(\omega)}$, cf. [34, Proposition 3.1] or [24, Theorem 3.3.2]. Here, $\mathbb{Z}^{(\omega)}$ denotes the direct sum of countably many copies of $\mathbb{Z}$, i.e., the subgroup of $\mathbb{Z}^{\omega}$ consisting of elements with finite support.
@@ 5. Uncountable groups
In what follows, given a cardinal $\kappa$, $2^{\kappa}$ denotes the topological space consisting of all functions $f:\kappa\rightarrow 2$, equipped with the product topology. In particular, basic clopen subsets of $2^{\kappa}$ are of the form $U_{\sigma}:=\{f\in 2^{\kappa}\mid\sigma\subseteq f\}$, where $\sigma$ is a finite partial function from $\kappa$ to $2$. Note that $2^{\kappa}$ is a profinite set. Our next result establishes item (1) of Theorem 1.3.
@@
for all $t\in B_{K}$ and extending linearly. The continuity of $\varphi$ follows from the continuity of $\varphi_{0}$, and our construction ensures that $\varphi(f)(t)\in\{0,1\}$ for all $f\in 2^{\kappa}$ and all $t\in B_{K}$. Moreover, $\varphi$ does not lift pointwise to a continuous $\psi:2^{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$, as pointwise restriction of such a $\psi$ to $F_{0}$ would induce a continuous pointwise extension $\psi_{0}:2^{\kappa}\rightarrow\mathrm{Hom}(F_{0},\mathbb{Z})$ of $\varphi_{0}$. We can therefore assume without loss of generality that $A$ is almost free.
@@
    for all $\eta<\kappa$ for which $\gamma_{\eta}\in E$, the set $A_{\gamma_{\eta+1}}/A_{\gamma}$ is not free.
@@ 6. Cohen forcing
Note that this establishes item (2) of Theorem 1.3. Before proving Theorem 6.1, we make the following observation.
@@
Thus, assume that $A$ is almost free, and hence $\kappa$ is regular. Since $A$ is nonfree, we can fix a $\subseteq$-increasing, continuous sequence $\vec{A}:=\langle A_{\alpha}\mid\alpha<\kappa\rangle$ such that
@@
    in every ccc forcing extension of $V^{\mathrm{Add}(\omega,\kappa)}$, there is no $\psi\in\mathrm{Hom}(F,\mathbb{Z})$ extending $\varphi$.
@@
We may thus assume that $\kappa>\aleph_{0}$ and $A$ is torsion-free. As in the proof of Theorem 5.1, we may assume that $A$ is almost free and $|F|=\kappa$. To see this, let $N$, $A_{0}$, $F_{0}$, $K_{0}$, $B_{F,0}$, and $B_{K,0}$ be as in the analogous argument in the earlier proof. Suppose that we are able to obtain, in $V^{\mathrm{Add}(\omega,\kappa)}$, a $\varphi_{0}\in\mathrm{Hom}(K_{0},\mathbb{Z})$ such that $\varphi_{0}[B_{K,0}]\subseteq\{0,1\}$ and $\varphi_{0}$ cannot be extended to a map $\psi_{0}\in\mathrm{Hom}(F_{0},\mathbb{Z})$ in any ccc forcing extension of $V^{\mathrm{Add}(\omega,\kappa)}$. Extend $\varphi_{0}$ to a map $\varphi\in\mathrm{Hom}(K,\mathbb{Z})$ by letting $\varphi(t)=0$ for all $t\in B_{K}\setminus B_{K,0}$. Then $\varphi\in V^{\mathrm{Add}(\omega,\kappa)}$. Moreover, if $\dot{\mathbb{P}}$ is an $\mathrm{Add}(\omega,\kappa)$-name for a ccc forcing notion, then $\varphi$ does not extend in $V^{\mathrm{Add}(\omega,\kappa)\ast\dot{\mathbb{P}}}$ to an element of $\mathrm{Hom}(F,\mathbb{Z})$, as the restriction of such an extension to $F_{0}$ would extend $\varphi_{0}$, contradicting our choice of $\varphi_{0}$.
@@
Note that $\mathrm{Add}(\omega,\kappa)\cong\mathrm{Add}(\omega,\kappa)\times\mathrm{Add}(\omega,\kappa)$, and hence we can assume that we are in fact working in an extension $V_{1}$ of the ground model $V$ by (the first copy of) $\mathrm{Add}(\omega,\kappa)$. In particular, by the inductive hypothesis, we can assume that, for every $\alpha\in S$, we have a $\varphi_{\alpha}\in\mathrm{Hom}(K_{\alpha}^{*},\mathbb{Z})$ such that
@@
  • in every ccc forcing extension of $V_{1}$, there is no $\psi_{\alpha}\in\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ extending $\varphi_{\alpha}$.
@@
Note that this construction continues to make sense in any outer model of $V_{1}$. Let $\dot{g}$ be the canonical $\mathrm{Add}(\omega,\kappa)$-name for the union of the generic filter, and note that $\dot{g}$ is a name for a function from $\kappa$ to $\{0,1\}$. Then let $\dot{\varphi}$ be an $\mathrm{Add}(\omega,\kappa)$-name for $\varphi(\dot{g})$. We claim that $\dot{\varphi}$ is forced to be as desired. To see this, let $\dot{\mathbb{P}}$ be an $\mathrm{Add}(\omega,\kappa)$-name for a ccc forcing notion, and suppose for the sake of contradiction that $\dot{\psi}$ is an $\mathrm{Add}(\omega,\kappa)\ast\dot{\mathbb{P}}$-name that is forced to be an element of $\mathrm{Hom}(F,\mathbb{Z})$ extending $\dot{\varphi}$.
@@
  • for all $\gamma<\alpha$, we have $\mathrm{dom}(p_{\gamma},q_{\gamma})\subseteq\alpha$.
@@
Since the map $p\mapsto\hat{p}$ is an automorphism of $\mathrm{Add}(\omega,\kappa)$, the filter $\hat{G}=\{\hat{p}\mid p\in G\}$ is generic over $V_{1}$, and the function $\hat{g}=\bigcup\hat{G}$ agrees with $g$ everywhere except $\alpha$, where it takes the opposite value. Without loss of generality, assume that $g(\alpha)=0$ and $\hat{g}(\alpha)=1$. For ease of comprehension, we then denote $(G,g)$ by $(G_{0},g_{0})$ and $(\hat{G},\hat{g})$ by $(G_{1},g_{1})$. Note also that $V[G_{0}]=V[G_{1}]$, and hence $H$ is $\mathbb{P}$-generic over both. For $i<2$, let $\varphi_{i}:=\varphi(g_{i})=\dot{\varphi}^{G_{i}}$, and let $\psi_{i}:=\dot{\psi}^{G_{i}\ast H}$.
By our assumptions about $\dot{\psi}$, we know that, for each $i<2$, $\psi_{i}\in\mathrm{Hom}(F,\mathbb{Z})$ extends $\varphi_{i}$. Moreover, by our choice of $\alpha$ and the fact that $g_{0}\restriction\alpha=g_{1}\restriction\alpha$, we know that $\psi_{0}\restriction F_{\alpha}=\psi_{1}\restriction F_{\alpha}$. Now define a homomorphism $\psi_{\alpha}\in\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ as follows. For all $z\in B^{*}_{F,\alpha}$, let $\psi_{\alpha}(z+F_{\alpha})=\psi_{1}(z)-\psi_{0}(z)$, and then extend linearly to the rest of $F_{\alpha}^{*}$. We claim that $\psi_{\alpha}$ extends $\varphi_{\alpha}$. The proof of this is exactly the same as that of the analogous fact in the proof of Theorem 5.1, so we leave it to the reader. But then, in $V_{1}[G\ast H]$, we have an element of $\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ extending $\varphi_{\alpha}$, contradicting our choice of $\varphi_{\alpha}$ and the fact that $V_{1}[G\ast H]$ is a ccc extension of $V_{1}$. ∎
@@ 7. Stone spaces
$N_{b}:=\{U\in S_{\kappa}\mid b\in U\}$
for $b\in\mathbb{B}_{\kappa}$.
@@
    Suppose now that $Y$ is a finite set and $\dot{f}$ is a $\mathbb{B}_{\kappa}$-name for a function from $X$ to $Y$. Define a function $\varphi_{\dot{f}}:S_{\kappa}\rightarrow Y^{X}$ as follows: for each $U\in S_{\kappa}$ and $x\in X$, let $\varphi_{\dot{f}}(U)(x)$ be the unique $y\in Y$ such that $\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\in U$. Note that such a $y$ must exist, since $\{\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\mid y\in Y\}$ is a finite maximal antichain in $\mathbb{B}_{\kappa}$ and therefore must intersect $U$.
It is routine to verify that $\dot{f}_{\varphi}$ as defined in Definition 7.1(1) is indeed a $\mathbb{B}_{\kappa}$ name for a function from $X$ to $Y$ and that $\varphi_{\dot{f}}$ as defined in Definition 7.1(2) is a continuous function from $S_{\kappa}$ to $Y^{X}$. (These facts will also follow from more general arguments appearing below in Section 8.) This translation, together with Theorem 6.3, allows us to replace the space $2^{\kappa}$ in Theorem 5.1 with the Stone space $S_{\kappa}$.
@@
We think of $\dot{\tau}$ as a $\mathbb{B}_{\kappa}$-name. Let $\dot{f}$ be the $\mathbb{B}_{\kappa}$-name for the restriction of $\tau$ to $B_{K}$; $\dot{f}$ is thus a name for a function from $B_{K}$ to $2$. Let $\varphi_{\dot{f}}$ be the continuous function from $S_{\kappa}$ to $2^{B_{K}}$ given Definition 7.1(2), and let $\varphi:S_{\kappa}\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ be defined by letting $\varphi(U)$ be the linear extension of $\varphi_{\dot{f}}(U)$ to all of $K$. Then $\varphi$ is also a continuous function.
We claim that $\varphi$ is as desired. We clearly have $\varphi(U)[B_{K}]\subseteq\{0,1\}$ for all $U\in S_{\kappa}$. Suppose for the sake of contradiction that $\psi:S_{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ is a continuous map and extends $\varphi$ pointwise. Let $\dot{f}_{\psi}$ be the $\mathbb{B}_{\kappa}$-name for a function from $F$ to $\mathbb{Z}$ given by Definition 7.1(1).
@@
Suppose for the sake of contradiction that there are $y,z\in F$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(y)+\dot{f}_{\psi}(z)\neq\dot{f}_{\psi}(y+z)$. By extending $b$, we may assume that there are $k_{y}$, $k_{z}$, and $k_{yz}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(y),\dot{f}_{\psi}(z),\dot{f}_{\psi}(y+z))=(k_{y},k_{z},k_{yz})$ and $k_{y}+k_{z}\neq k_{yz}$. Let $U\in S_{\kappa}$ be such that $b\in U$. Then by the definition of $\dot{f}_{\psi}$, we have $\psi(U)(y)=k_{y}$, $\psi(U)(z)=k_{z}$, and $\psi(U)(y+z)=k_{yz}$, contradicting the fact that $\psi(U)\in\mathrm{Hom}(F,\mathbb{Z})$. ∎
@@
Suppose for the sake of contradiction that there is $t\in B_{K}$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(t)\neq\dot{\tau}(t)$. By extending $b$ if necessary, we may assume without loss of generality that there are $k_{\psi},k_{\tau}\in\mathbb{Z}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(t),\dot{\tau}(t))=(k_{\psi},k_{\tau})$. Let $U\in S_{\kappa}$ be such that $b\in U$. Then by the definition of $\varphi_{\dot{f}}$, we have $\varphi_{\dot{f}}(U)(t)=\varphi(U)(t)=k_{\psi}$, and by the definition of $\dot{f}_{\psi}$, we have $\psi(U)(t)=k_{\tau}$, contradicting the fact that $\psi$ extends $\varphi$ pointwise. ∎

### v5

@@ 4. Countable groups
We can thus injectively enumerate $B_{K}$ as $\langle t_{n}\mid n<\omega\rangle$ and, for each $n<\omega$, let $K_{n}$ denote the subgroup of $K$ generated by $\{t_{k}\mid k<n\}$. We will define our homomorphism $\varphi:K\rightarrow\mathbb{Z}$ by recursion, requiring that $\varphi(t_{n})\in\{0,1\}$ for all $n<\omega$. We will do this by constructing a strictly increasing sequence $\langle n_{i}\mid i<\omega\rangle$ of natural numbers and, for each $i<\omega$, a homomorphism $\varphi_{i}:K_{n_{i}}\rightarrow\mathbb{Z}$ in such a way that $\varphi_{i}[\{t_{k}\mid k<n_{i}\}]\subseteq\{0,1\}$ and, for all $i<j<\omega$, $\varphi_{j}\restriction K_{n_{i}}=\varphi_{i}$. In the process, we will also define a $\subseteq$-increasing sequence $\langle Y_{n}\mid n<\omega\rangle$ of finite subsets of $B_{F}$ such that, for all $i<\omega$, letting $F_{i}$ denote the subgroup of $F$ generated by $Y_{i}$, we have $K_{n_{i}}\subseteq F_{i}$.
@@
Then $s_{i}^{*}\in K$ is of the form $j_{i}z_{i}-r$, where $r\in F_{i}$. Moreover, since all of the coefficients $b_{k}$ for $k\in[n_{i},n_{i+1})$ are divisible by $j_{i}$, it follows that $s_{i}^{*}$ is divisible by $j_{i}$ (in $K$). Since $r=j_{i}z_{i}-s_{i}^{*}$ is in $F_{i}$ and $F_{i}$ is generated by a subset of $B_{F}$, this means that there exists $r^{*}\in F_{i}$ such that $r=j_{i}r^{*}$. We get $s_{i}^{*}=j_{i}(z_{i}-r^{*})\in K$ which implies $z_{i}-r^{*}\in K$ using the fact that $A$ is torsion-free. In particular, $\pi(z_{i})=\pi(r^{*})$, so $\pi(z_{i})\in\pi[F_{i}]$, contradicting the fact that $\pi(z_{i})=x_{i}\notin\pi[F_{i}]$. ∎
@@
In the formulation of Theorem 4.2, we could not have hoped to find $\varphi$ with $(\varphi\restriction B_{K})^{-1}\{1\}$ finite, let alone a $\varphi$ satisfying the stronger condition (2) from Proposition 4.4. The reason is that any given countable torsion-free group $A$, being the union of a countable chain of finitely generated free subgroups, is isomorphic to a (pure) subgroup of $\mathbb{Z}^{\omega}/\mathbb{Z}^{(\omega)}$, cf. [26, Proposition 3.1] or [25, Theorem 3.3.2]. Here, $\mathbb{Z}^{(\omega)}$ denotes the direct sum of countably many copies of $\mathbb{Z}$, i.e., the subgroup of $\mathbb{Z}^{\omega}$ consisting of elements with finite support.
@@ 5. Uncountable groups
In what follows, given a cardinal $\kappa$, $2^{\kappa}$ denotes the topological space consisting of all functions $f:\kappa\rightarrow 2$, equipped with the product topology. In particular, basic clopen subsets of $2^{\kappa}$ are of the form $U_{\sigma}:=\{f\in 2^{\kappa}\mid\sigma\subseteq f\}$, where $\sigma$ is a finite partial function from $\kappa$ to $2$. Note that $2^{\kappa}$ is a profinite set.
@@
for all $t\in B_{K}$ and extending linearly. The continuity of $\varphi$ follows from the continuity of $\varphi_{0}$, and our construction ensures that $\varphi(f)(t)\in\{0,1\}$ for all $f\in 2^{\kappa}$ and all $t\in B_{K}$. Moreover, $\varphi$ does not lift pointwise to a continuous $\psi:2^{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$, as pointwise restriction of such a $\psi$ to $F_{0}$ would induce a continuous pointwise extension $\psi_{0}:2^{\kappa}\rightarrow\mathrm{Hom}(F_{0},\mathbb{Z})$ of $\varphi_{0}$.
@@
    for all $\eta<\kappa$ for which $\gamma_{\eta}\in E$, the set $A_{\gamma_{\eta+1}}/A_{\gamma_{\eta}}$ is not free.
@@ 6. Cohen forcing
Note that this establishes clause (1) of Theorem 1.3. Before proving Theorem 6.1, we make the following observation.
@@
Thus, assume that $A$ is almost free, and hence, by Shelah’s aforementioned singular compactness theorem, $\kappa$ is regular. Since $A$ is nonfree, we can fix a $\subseteq$-increasing, continuous sequence $\vec{A}:=\langle A_{\alpha}\mid\alpha<\kappa\rangle$ such that
@@
    if $\mathbb{Q}\in V$ is a ccc forcing notion, then, in $V^{\mathrm{Add}(\omega,\kappa)\times\mathbb{Q}}$, there is no $\psi\in\mathrm{Hom}(F,\mathbb{Z})$ extending $\varphi$.
@@
We may thus assume that $\kappa>\aleph_{0}$ and $A$ is torsion-free. As in the proof of Theorem 5.1, we may assume that $A$ is almost free and $|F|=\kappa$. To see this, let $N$, $A_{0}$, $F_{0}$, $K_{0}$, $B_{F,0}$, and $B_{K,0}$ be as in the analogous argument in the earlier proof. Suppose that we are able to obtain, in $V^{\mathrm{Add}(\omega,\kappa)}$, a $\varphi_{0}\in\mathrm{Hom}(K_{0},\mathbb{Z})$ such that $\varphi_{0}[B_{K,0}]\subseteq\{0,1\}$ and, for every ccc $\mathbb{Q}\in V$, $\varphi_{0}$ cannot be extended to a map $\psi_{0}\in\mathrm{Hom}(F_{0},\mathbb{Z})$ in $V^{\mathrm{Add}(\omega,\kappa)\times\mathbb{Q}}$. Extend $\varphi_{0}$ to a map $\varphi\in\mathrm{Hom}(K,\mathbb{Z})$ by letting $\varphi(t)=0$ for all $t\in B_{K}\setminus B_{K,0}$. Then $\varphi\in V^{\mathrm{Add}(\omega,\kappa)}$. Moreover, if $\mathbb{Q}\in V$ is a ccc forcing notion, then $\varphi$ does not extend in $V^{\mathrm{Add}(\omega,\kappa)\times\mathbb{Q}}$ to an element of $\mathrm{Hom}(F,\mathbb{Z})$, as the restriction of such an extension to $F_{0}$ would extend $\varphi_{0}$, contradicting our choice of $\varphi_{0}$.
@@
Note that $\mathrm{Add}(\omega,\kappa)\cong\mathrm{Add}(\omega,\kappa)\times\mathrm{Add}(\omega,\kappa)$; move to an extension $V_{1}$ of the ground model $V$ by (the first copy of) $\mathrm{Add}(\omega,\kappa)$. In particular, by the inductive hypothesis, we can assume that, for every $\alpha\in S$, we have a $\varphi_{\alpha}\in\mathrm{Hom}(K_{\alpha}^{*},\mathbb{Z})$ such that
@@
  • for every ccc $\mathbb{Q}\in V$, in $V_{1}^{\mathbb{Q}}$ there is no $\psi_{\alpha}\in\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ extending $\varphi_{\alpha}$.
@@
Note that this construction continues to make sense in any outer model of $V_{1}$. Let $\dot{g}$ be the canonical $\mathrm{Add}(\omega,\kappa)$-name for the union of the generic filter, and note that $\dot{g}$ is a name for a function from $\kappa$ to $\{0,1\}$. Then let $\dot{\varphi}$ be an $\mathrm{Add}(\omega,\kappa)$-name for $\varphi(\dot{g})$. We claim that $\dot{\varphi}$ is forced to be as desired. To see this, let $\mathbb{Q}\in V$ be a ccc forcing notion, and suppose for the sake of contradiction that $\dot{\psi}\in V_{1}$ is an $\mathrm{Add}(\omega,\kappa)\times\mathbb{Q}$-name that is forced to be an element of $\mathrm{Hom}(F,\mathbb{Z})$ extending $\dot{\varphi}$.
@@
  • for all $\gamma<\alpha$, we have $\mathrm{dom}(p_{\gamma})\subseteq\alpha$.
@@
[…]
By our assumptions about $\dot{\psi}$, we know that, for each $i<2$, $\psi_{i}\in\mathrm{Hom}(F,\mathbb{Z})$ extends $\varphi_{i}$. Moreover, by our choice of $\alpha$ and the fact that $g_{0}\restriction\alpha=g_{1}\restriction\alpha$, we know that $\psi_{0}\restriction F_{\alpha}=\psi_{1}\restriction F_{\alpha}$. Now define a homomorphism $\psi_{\alpha}\in\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ as follows. For all $z\in B^{*}_{F,\alpha}$, let $\psi_{\alpha}(z+F_{\alpha})=\psi_{1}(z)-\psi_{0}(z)$, and then extend linearly to the rest of $F_{\alpha}^{*}$. We claim that $\psi_{\alpha}$ extends $\varphi_{\alpha}$. The proof of this is exactly the same as that of the analogous fact in the proof of Theorem 5.1, so we leave it to the reader. But then, in $V_{1}[G\times H]$, we have an element of $\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ extending $\varphi_{\alpha}$, contradicting our choice of $\varphi_{\alpha}$ and the fact that $V_{1}[G\times H]$ is an extension of $V_{1}$ by $\mathrm{Add}(\omega,\kappa)\times\mathbb{Q}$, which is a ccc forcing notion in $V$. ∎
@@ 7. Stone spaces
$N_{b}:=\{\mathcal{U}\in S_{\kappa}\mid b\in\mathcal{U}\}$
for $b\in\mathbb{B}^{+}_{\kappa}$.
@@
    Suppose now that $Y$ is a finite set and $\dot{f}$ is a $\mathbb{B}_{\kappa}$-name for a function from $X$ to $Y$. Define a function $\varphi_{\dot{f}}:S_{\kappa}\rightarrow Y^{X}$ as follows: for each $\mathcal{U}\in S_{\kappa}$ and $x\in X$, let $\varphi_{\dot{f}}(\mathcal{U})(x)$ be the unique $y\in Y$ such that $\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\in\mathcal{U}$. Note that such a $y$ must exist, since $\{\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\mid y\in Y\}$ is a finite maximal antichain in $\mathbb{B}_{\kappa}$ and therefore must intersect $\mathcal{U}$.
It is routine to verify that $\dot{f}_{\varphi}$ as defined in Definition 7.1(1) is indeed a $\mathbb{B}_{\kappa}$-name for a function from $X$ to $Y$ and that $\varphi_{\dot{f}}$ as defined in Definition 7.1(2) is a continuous function from $S_{\kappa}$ to $Y^{X}$. (These facts will also follow from more general arguments appearing below in Section 8.) This translation, together with Theorem 6.3, allows us to replace the space $2^{\kappa}$ in Theorem 5.1 with the Stone space $S_{\kappa}$.
@@
We think of $\dot{\tau}$ as a $\mathbb{B}_{\kappa}$-name. Let $\dot{f}$ be the $\mathbb{B}_{\kappa}$-name for the restriction of $\tau$ to $B_{K}$; $\dot{f}$ is thus a name for a function from $B_{K}$ to $2$. Let $\varphi_{\dot{f}}$ be the continuous function from $S_{\kappa}$ to $2^{B_{K}}$ given by Definition 7.1(2), and let $\varphi:S_{\kappa}\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ be defined by letting $\varphi(\mathcal{U})$ be the linear extension of $\varphi_{\dot{f}}(\mathcal{U})$ to all of $K$. Then $\varphi$ is also a continuous function.
We claim that $\varphi$ is as desired. We clearly have $\varphi(\mathcal{U})[B_{K}]\subseteq\{0,1\}$ for all $\mathcal{U}\in S_{\kappa}$. Suppose for the sake of contradiction that $\psi:S_{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ is a continuous map and extends $\varphi$ pointwise. Let $\dot{f}_{\psi}$ be the $\mathbb{B}_{\kappa}$-name for a function from $F$ to $\mathbb{Z}$ given by Definition 7.1(1).
@@
Suppose for the sake of contradiction that there are $y,z\in F$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(y)+\dot{f}_{\psi}(z)\neq\dot{f}_{\psi}(y+z)$. By extending $b$, we may assume that there are $k_{y}$, $k_{z}$, and $k_{yz}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(y),\dot{f}_{\psi}(z),\dot{f}_{\psi}(y+z))=(k_{y},k_{z},k_{yz})$ and $k_{y}+k_{z}\neq k_{yz}$. Let $\mathcal{U}\in S_{\kappa}$ be such that $b\in\mathcal{U}$. Then by the definition of $\dot{f}_{\psi}$, we have $\psi(\mathcal{U})(y)=k_{y}$, $\psi(\mathcal{U})(z)=k_{z}$, and $\psi(\mathcal{U})(y+z)=k_{yz}$, contradicting the fact that $\psi(\mathcal{U})\in\mathrm{Hom}(F,\mathbb{Z})$. ∎
@@
Suppose for the sake of contradiction that there is $t\in B_{K}$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(t)\neq\dot{\tau}(t)$. By extending $b$ if necessary, we may assume without loss of generality that there are $k_{\psi},k_{\tau}\in\mathbb{Z}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(t),\dot{\tau}(t))=(k_{\psi},k_{\tau})$. Let $\mathcal{U}\in S_{\kappa}$ be such that $b\in\mathcal{U}$. Then by the definition of $\varphi_{\dot{f}}$, we have $\varphi_{\dot{f}}(\mathcal{U})(t)=\varphi(\mathcal{U})(t)=k_{\psi}$, and by the definition of $\dot{f}_{\psi}$, we have $\psi(\mathcal{U})(t)=k_{\tau}$, contradicting the fact that $\psi$ extends $\varphi$ pointwise. ∎